OpenStudy (anonymous):

evaluate the limit algebraically... lim x-->2 (x-2)/(sqrt(x)-sqrt(4-x))

OpenStudy (anonymous):

\[\lim_{x\rightarrow 2}\frac{x-2}{\sqrt{x}-\sqrt{4-x}}\] like that?

OpenStudy (anonymous):

first make sure that, if you replace x by 2, you get 0/0, which you do. then multiply numerator and denominator by the conjugate of the denominator, i.e. multiply by \[\frac{\sqrt{x}+\sqrt{4-x}}{\sqrt{x}+\sqrt{4-x}}\]

OpenStudy (anonymous):

your new denominator will be \[2x-4=2(x-2)\] so you can cancel with the \[x-2\] in the numerator, leaving only \[\frac{\sqrt{x}+\sqrt{4-x}}{2}\] and now replace x by 2

OpenStudy (anonymous):

and don't kick out farmdawg, he doesn't like it

OpenStudy (anonymous):

haha what about a farmdawg? other than that thank you...